OpenStudy (anonymous):

If the average of four numbers is 22, and three of the numbers are 7, 11, and 18, then what is the fourth number?

OpenStudy (anonymous):

52 because 52+11+7+18= 88 and 88 divided by 4 = 22

OpenStudy (mathmale):

average = (sum of all the numbers) / (count of the numbers...how many there are). Here we know the average is 22, and we know 3 of the 4 numbers, so we represent the fourth number by x. Add up the four numbers. Sum of the numbers is: ???? How many numbers are there? ???? Divde the sum by the count. Equate the result to the average (22). How would you now solve for x?

OpenStudy (mathmale):

Isaiah, why the x in the denom.?

OpenStudy (mathmale):

Isn't the average = to 22?

OpenStudy (anonymous):

\[\frac{1}{4} (x+18+11+7)=22 \]x = 52

OpenStudy (mathmale):

Nice work, Robert. But you might want to lead Megan through the steps involved here.

OpenStudy (jagatuba):

\[\frac{ 7+11+18+x }{ 4 }=22\]

OpenStudy (mathmale):

Again, right on. But how do you solve this equation for x? Megan may want to know.

OpenStudy (mathmale):

(For jagatuba only: Thanks for referring me to Preetha a week or so ago. I did contact her and find her to be very helpful!)

OpenStudy (anonymous):

@megan177 The average of n numbers is the sum of the n numbers divided by n.

OpenStudy (jagatuba):

\[36+x=22\times4\] \[x=88-36\]

OpenStudy (mathmale):

Right on. But what about the algebraic steps involved in solving that equation for x?

OpenStudy (mathmale):

Cool, Jag! Megan, how are you doing?

OpenStudy (jagatuba):

@mathmale YW. Glad she could help.

OpenStudy (mathmale):

Thanks. Megan? Still with us?

OpenStudy (mathmale):

@megan177 ?

OpenStudy (anonymous):

Megan is "off line"

OpenStudy (anonymous):

relax I was off line, my computer froze. okay I get it now, thanks guys